# Evaluating Diversity in Automatic Poetry Generation

### v1

@@ Appendix A Appendix
Model | EN | DE |
GPT2-small | 124M | 124M |
GPT2-large | 774M | 774M |
GPTneo-small | 125M | - |
GPTneo-xl | 1.3B | - |
byGPT5-base | 140M | 140M |
byGPT5-medium | 290M | 290M |
@@
A.1 Rhyme
DE | EN | |||||
Model | Entropy | KL Divergence | Cross Entropy | Entropy | KL Divergence | Cross Entropy |
HUMAN | 2.90 | 0.00 | 2.90 | 3.10 | 0.00 | 3.10 |
DeepSpeare | 2.97 | 0.55 | 3.52 | 3.16 | 0.48 | 3.64 |
SA | 3.14 | 1.43 | 4.57 | 3.22 | 1.17 | 4.39 |
ByGPT5 | 2.89 | 1.23 | 4.12 | 2.92 | 1.08 | 4.00 |
ByGPT5 | 3.13 | 1.09 | 4.21 | 2.91 | 1.13 | 4.04 |
GPT2 | 2.86 | 1.26 | 4.13 | 2.97 | 1.06 | 4.03 |
GPT2 | 3.16 | 1.13 | 4.29 | 2.99 | 1.03 | 4.03 |
GPTNeo | - | - | - | 2.80 | 1.18 | 3.99 |
GPTNeo | - | - | - | 3.16 | 0.96 | 4.11 |
LLaMA2 | 2.93 | 1.18 | 4.11 | 3.24 | 0.71 | 3.95 |
LLaMA2 | 3.18 | 1.04 | 4.21 | 3.24 | 0.71 | 3.95 |
LLaMA3 | 3.27 | 0.83 | 4.10 | 3.45 | 0.56 | 4.02 |
ByGPT5 | 3.17 | 0.67 | 3.84 | 3.22 | 0.83 | 4.05 |
ByGPT5 | 3.16 | 0.58 | 3.75 | 3.38 | 0.54 | 3.92 |
GPT2 | 2.98 | 0.99 | 3.97 | 3.41 | 0.61 | 4.02 |
GPT2 | 3.11 | 1.04 | 4.14 | 3.22 | 0.85 | 4.07 |
GPTNeo | - | - | - | 3.43 | 0.45 | 3.89 |
GPTNeo | - | - | - | 3.29 | 0.83 | 4.11 |
LLaMA2 | 2.69 | 1.33 | 4.02 | 2.89 | 0.95 | 3.85 |
LLaMA2 | 3.11 | 0.71 | 3.83 | 2.67 | 1.07 | 3.74 |
LLaMA3 | 2.98 | 1.06 | 4.04 | 2.58 | 0.94 | 3.52 |
Appendix B Decoding methods
Method | GPT2-large | p-GPTneo-xl | p-byGPT5-medium |
0.002 | 0.89 | 0.006 | |
0.000 | 0.96 | 0.006 | |
0.000 | 0.98 | 0.004 | |
0.002 | 0.99 | 0.010 | |
0.004 | 0.98 | 0.004 | |
0.012 | 0.99 | 0.020 | |
0.002 | 0.99 | 0.010 | |
0.000 | 0.99 | 0.008 | |
0.002 | 0.98 | 0.004 | |
0.004 | 0.99 | 0.012 | |
0.004 | 0.94 | 0.004 |
Method | mean | std | min | max | |
0.002 | 0.55 | 0.06 | 0.33 | 0.77 | |
0.000 | 0.57 | 0.06 | 0.37 | 0.82 | |
0.000 | 0.58 | 0.06 | 0.39 | 0.72 | |
0.002 | 0.57 | 0.07 | 0.38 | 0.81 | |
0.004 | 0.57 | 0.06 | 0.37 | 0.97 | |
0.012 | 0.59 | 0.07 | 0.44 | 0.99 | |
0.002 | 0.56 | 0.06 | 0.37 | 0.91 | |
0.000 | 0.58 | 0.07 | 0.33 | 0.84 | |
0.002 | 0.57 | 0.07 | 0.42 | 0.95 | |
0.004 | 0.57 | 0.07 | 0.39 | 0.98 | |
0.004 | 0.56 | 0.06 | 0.34 | 0.94 |
Method | mean | std | min | max | h | |
29.83 | 13.64 | 4 | 150 | 0.75 | 5.46 | |
37.77 | 20.58 | 4 | 150 | 0.87 | 4.48 | |
35.67 | 8.85 | 13 | 64 | 0.86 | 1.89 |
Model | lang | real rhymes | repetitions |
en | 0.72 | 0.00 | |
en | 0.84 | 0.00 | |
en | 0.44 | 0.08 | |
en | 0.53 | 0.21 | |
de | 0.57 | 0.00 | |
de | 0.66 | 0.00 |
method | real rhymes | repetitions |
0.156 | 0.022 | |
0.196 | 0.052 | |
0.164 | 0.071 | |
0.176 | 0.220 | |
0.170 | 0.103 | |
0.181 | 0.124 | |
0.146 | 0.030 | |
0.168 | 0.131 | |
0.191 | 0.112 | |
0.175 | 0.153 | |
0.194 | 0.087 |
Method | ATTR | MATTR | MTLD |
0.804 | 0.794 | 64.679 | |
0.757 | 0.744 | 44.942 | |
0.716 | 0.717 | 36.561 | |
0.537 | 0.537 | 17.297 | |
0.720 | 0.690 | 33.151 | |
0.419 | 0.461 | 14.730 | |
0.586 | 0.622 | 21.757 | |
0.611 | 0.588 | 21.005 | |
0.766 | 0.754 | 46.107 | |
0.642 | 0.637 | 24.289 | |
0.797 | 0.779 | 58.490 |

### v2

@@ Appendix A Appendix
A.1 DeepSpeare and SA
Deepspeare Lau et al. (2018) is specifically designed for poetry generation. Its core architecture consists of an LSTM language model, a pentameter model (specifically designed to learn iambic meter) and a rhyme model. During training, it takes sonnets as input data (three quatrains followed by a couplet) but ultimately processes the contained quatrains by splitting any given sonnet. The rhyme model processes ending words of quatrain verses and uses a margin-based loss to discriminate between rhyming and non-rhyming words. It is not limited to specific rhyme patterns but assumes that rhymes exist in the data. At inference time, Deepspeare generates quatrains.
Structured Adversary.
Like Deepspeare, Structured Adversary (SA) Jhamtani et al. (2019) incorporates different components: an LSTM language model and a discriminator used to decide whether line endings are typical for poetry. Both components are organized in an adversarial setup, where the language model acts as a generator, trying to generate poems that are misclassified by the discriminator, while the discriminator is trained to distinguish generated poems from real ones. SA is trained with sonnets as input data. At inference time, it generates quatrains.
A.2 Training
DeepSpeare
DeepSpeare (Lau et al., 2018) leverages pretrained static word vectors. We use QuaTrain and SonNet to train our own Word2vec embeddings (Mikolov et al., 2013) and the final sonnet models respectively. For the sonnet model training, we use a batch size of 128 and apply early stopping with a patience of 5 epochs; default settings are maintained for the other hyperparameters.
SA
Style-un/conditioned LLMs
We train all LLMs for 50 epochs on our train set using the paged AdamW optimizer with a weight decay of 0.001, a learning rate of 4e-05, a cosine learning rate decay with a 3% warmup ratio, and early stopping with patience of 5 epochs. As we run experiments on GPUs with varying memory capacities ranging from 12GB to 80GB, and with models that drastically differ in size, to achieve as much consistency as possible, we either train models with a batch size of 128 or accumulate the batches to reach a size of 128. For LLaMA, we use 4-bit quantization and LORA (Hu et al., 2021); the corresponding parameters are list below:
target modules: q_proj, v_proj, k_proj, o_proj, embedded_tokens
lora alpha: 16
lora dropout: 0.05
r: 16
A.3 Evaluation Results
@@
Length
Rhyme
Table 8 shows the entropy of the rhyme distributions in each sample as well as the distances of the distributions to that in the human data, measured by KL divergence. Figure 3 demonstrates the human rhyme distribution as well as the best, worst, and an average fit distributions in terms of KL divergence. Figures 5, 6/7, and 8/9 demonstrate the rhyme distributions for the poetry specific models, unconditioned and conditioned LLMs, respectively.
| DE | EN | |||
| Model | Entropy | KL Divergence | Entropy | KL Divergence |
| HUMAN | 2.90 | 0.00 | 3.10 | 0.00 |
| DeepSpeare | 2.97 | 0.55 | 3.16 | 0.48 |
| SA | 3.14 | 1.43 | 3.22 | 1.17 |
| ByGPT5 | 2.89 | 1.23 | 2.92 | 1.08 |
| ByGPT5 | 3.13 | 1.09 | 2.91 | 1.13 |
| GPT2 | 2.86 | 1.26 | 2.97 | 1.06 |
| GPT2 | 3.16 | 1.13 | 2.99 | 1.03 |
| GPTNeo | - | - | 2.80 | 1.18 |
| GPTNeo | - | - | 3.16 | 0.96 |
| LLaMA2 | 2.93 | 1.18 | 3.24 | 0.71 |
| LLaMA2 | 3.18 | 1.04 | 3.24 | 0.71 |
| LLaMA3 | 3.27 | 0.83 | 3.45 | 0.56 |
| ByGPT5 | 3.17 | 0.67 | 3.22 | 0.83 |
| ByGPT5 | 3.16 | 0.58 | 3.38 | 0.54 |
| GPT2 | 2.98 | 0.99 | 3.41 | 0.61 |
| GPT2 | 3.11 | 1.04 | 3.22 | 0.85 |
| GPTNeo | - | - | 3.43 | 0.45 |
| GPTNeo | - | - | 3.29 | 0.83 |
| LLaMA2 | 2.69 | 1.33 | 2.89 | 0.95 |
| LLaMA2 | 3.11 | 0.71 | 2.67 | 1.07 |
| LLaMA3 | 2.98 | 1.06 | 2.58 | 0.94 |
Best model
We rank the models for each dimension and then average the ranks across the five dimensions to determine the overall rankings. For dimensions with multiple metrics, such as the three memorization metrics (due to different evaluation levels) and the three lexical metrics (measuring local or global lexical diversity), we first rank the models according to each metric and then average these ranks to represent that dimension. For dimensions primarily based on distributions, we use metrics that measure the distance/similarity of their distributions from human data: KL divergence for rhyme and histogram intersection for length. The results are shown in Table 9 and 10 for German and English respectively.
| Language | Model | Size | Conditioned | semantic | lexical | length | rhyme | memorization | avg_rank |
| de | BYGPT5 | L | TRUE | 2.0 | 4.0 | 5.0 | 3.0 | 1.7 | 3.1 |
| de | BYGPT5 | S | TRUE | 3.5 | 6.0 | 4.0 | 2.0 | 1.3 | 3.4 |
| de | SA | - | - | 1.0 | 2.7 | 1.0 | 16.0 | 2.0 | 4.5 |
| de | DS | - | - | 5.0 | 10.3 | 12.0 | 1.0 | 1.0 | 5.9 |
| de | BYGPT5 | S | FALSE | 6.0 | 11.0 | 2.0 | 10.0 | 2.7 | 6.3 |
| de | BYGPT5 | L | FALSE | 4.0 | 8.3 | 6.0 | 13.0 | 3.0 | 6.9 |
| de | LLAMA3 | - | FALSE | 9.5 | 6.3 | 9.0 | 5.0 | 6.0 | 7.2 |
| de | LLAMA3 | - | TRUE | 6.5 | 7.3 | 8.0 | 9.0 | 5.7 | 7.3 |
| de | LLAMA2 | S | TRUE | 13.5 | 13.0 | 3.0 | 4.0 | 4.0 | 7.5 |
| de | GPT2 | L | TRUE | 12.5 | 4.7 | 7.0 | 6.0 | 8.3 | 7.7 |
| de | LLAMA2 | L | FALSE | 9.5 | 2.7 | 16.0 | 12.0 | 5.3 | 9.1 |
| de | LLAMA2 | S | FALSE | 8.0 | 10.0 | 15.0 | 8.0 | 5.0 | 9.2 |
| de | GPT2 | L | FALSE | 14.0 | 5.7 | 10.0 | 14.0 | 8.7 | 10.5 |
| de | GPT2 | S | TRUE | 15.0 | 15.0 | 11.0 | 7.0 | 6.3 | 10.9 |
| de | LLAMA2 | L | TRUE | 12.5 | 13.0 | 13.0 | 15.0 | 8.0 | 12.3 |
| de | GPT2 | S | FALSE | 13.5 | 16.0 | 14.0 | 11.0 | 7.7 | 12.4 |
| Language | Model | Size | Conditioned | semantic | lexical | length | rhyme | memorization | avg_rank |
| en | BYGPT5 | S | TRUE | 3.5 | 11.7 | 4.0 | 3.0 | 2.0 | 4.8 |
| en | SA | - | - | 1.0 | 4.0 | 1.0 | 19.0 | 1.0 | 5.2 |
| en | BYGPT5 | L | TRUE | 2.0 | 9.7 | 5.0 | 9.0 | 1.7 | 5.5 |
| en | DS | - | - | 3.5 | 9.0 | 17.0 | 2.0 | 2.3 | 6.8 |
| en | LLAMA2 | S | FALSE | 17.5 | 5.7 | 2.0 | 6.0 | 4.7 | 7.2 |
| en | LLAMA3 | - | TRUE | 12.0 | 1.7 | 9.0 | 11.0 | 3.3 | 7.4 |
| en | GPT2 | L | TRUE | 9.0 | 9.0 | 6.0 | 5.0 | 9.3 | 7.7 |
| en | LLAMA2 | L | TRUE | 12.0 | 5.0 | 7.0 | 12.0 | 4.0 | 8.0 |
| en | LLAMA2 | S | TRUE | 7.0 | 3.3 | 13.0 | 16.0 | 1.3 | 8.1 |
| en | LLAMA3 | - | FALSE | 13.0 | 3.0 | 16.0 | 4.0 | 9.0 | 9.0 |
| en | LLAMA2 | L | FALSE | 9.0 | 6.3 | 15.0 | 7.0 | 10.3 | 9.5 |
| en | GPT2 | S | TRUE | 17.5 | 14.0 | 3.0 | 10.0 | 3.7 | 9.6 |
| en | BYGPT5 | L | FALSE | 5.5 | 15.7 | 10.0 | 17.0 | 3.0 | 10.2 |
| en | BYGPT5 | S | FALSE | 5.5 | 17.3 | 8.0 | 18.0 | 2.7 | 10.3 |
| en | GPTNEO | L | TRUE | 13.5 | 13.0 | 19.0 | 1.0 | 10.0 | 11.3 |
| en | GPTNEO | S | TRUE | 16.0 | 17.0 | 11.0 | 8.0 | 5.7 | 11.5 |
| en | GPT2 | L | FALSE | 10.5 | 11.0 | 12.0 | 15.0 | 11.3 | 12.0 |
| en | GPT2 | S | FALSE | 17.0 | 19.0 | 14.0 | 14.0 | 11.7 | 15.1 |
| en | GPTNEO | S | FALSE | 17.5 | 20.0 | 18.0 | 13.0 | 12.0 | 16.1 |
| en | GPTNEO | L | FALSE | 17.5 | 14.7 | 20.0 | 20.0 | 11.3 | 16.7 |
| DE | EN | |||||||||
| System | HUMAN | overall | semantic | lexical | rhyme | HUMAN | overall | semantic | lexical | rhyme |
| (ByGPT5) | (SA) | (SA) | (DeepSpeare) | (ByGPT5) | (SA) | (LLaMA3) | (GPTNeo) | |||
| Best | 12 | 2 | 0 | - | 1 | 3 | 1 | 0 | 6 | 5 |
| Worst | 0 | 1 | 8 | - | 6 | 2 | 2 | 11 | 0 | 0 |
| BWS | 0.8 | 0.07 | -0.53 | - | -0.33 | 0.07 | -0.07 | -0.73 | 0.4 | 0.33 |
| Quatrain | System | ||||
ByGPT5 | ||||
ByGPT5 | ||||
DeepSpeare | ||||
HUMAN | ||||
HUMAN |
| Quatrain | System | ||||
LLaMA3 | ||||
GPTNeo | ||||
LLaMA3 | ||||
HUMAN | ||||
HUMAN |